The ratio measure of two angles that are complementary is 4:5. What is the measure of the two angles?
Answer by checkley77(12569)

4+5=9
90/9=10
4*10=40 DEGREES ANS.
5*10=50 DEGREES ANS.
PROOF40+50=90
4:5=40:50